OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve: sin(x+y)=6x-6y at pts (pi,pi)

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

well I know the derivative of it: (-(6-cos(x+y))/(6+cos(x+y)) but I can't figure out the equation, or maybe my homework program won't take it

OpenStudy (anonymous):

so my derivative was the same just plug in pi and youll get cos2pi which is one and simplifies easy

OpenStudy (anonymous):

but i got a 6-cos(x+y) on top and when i plugged pi in i got 5/7

OpenStudy (anonymous):

right, but it wants it in a tangent line equation so I don't know how to put it in that form...

OpenStudy (anonymous):

does it ask for a point slope form?

OpenStudy (anonymous):

no it asks for a tangent line equation, which I am unsure what that means..

OpenStudy (anonymous):

a tangent line equation usually refers to the derivative or dy/dx for implicit differentiation so maybe try without your negative in front of your derivative

OpenStudy (anonymous):

still says incorrect when my line is: (5/7)x+(pi-5/7)=0

OpenStudy (anonymous):

5/7 is the slope so what you want if youre finding a tangent line is most likely a point slope or, y-y1=m(x-x1) where you plug in pi and pi for for x1 and y1 and 5/7 for slope so youll get (y-pi)=5/7(x-pi)

OpenStudy (anonymous):

oh man that worked, and makes more sense! thank you!

OpenStudy (anonymous):

anytime they ask for an equation 90% of the time its point slope where x1 and y1 is your point given you seem to know the rest good luck!

OpenStudy (anonymous):

I am also in calc this year and even i can say it's quite challenging you have make sure you know why youre doing it not just what youre doing hope it goes well